### v2

# Randomized SINDy

Abstract.
This paper presents “randomized SINDy”, a sequential machine learning algorithm designed for dynamic data that has a time-dependent structure. It employs a probabilistic approach, with its PAC learning property rigorously proven through the mathematical theory of functional analysis. The algorithm dynamically predicts using a learned probability distribution of predictors, updating weights via gradient descent and a proximal algorithm to maintain a valid probability density. Inspired by SINDy ([2]), it incorporates feature augmentation and Tikhonov regularization. For multivariate normal weights, the proximal step is omitted to focus on parameter estimation. The algorithm’s effectiveness is demonstrated through experimental results in regression and binary classification using real-world data.
1. Introduction
Sequential machine learning represents a family of efficient and scalable algorithms to incrementally build predictive models from a sequence of data ([14]). Traditional machine learning algorithms have a high computational cost for retraining in the presence of new data while sequential algorithms are suitable for applications with a high volume of data sampled dynamically and quickly. Huge databases that grow dynamically over time require scalable machine learning algorithms [9, 10, 11].
Traditional machine learning algorithms assume that the data is independent and identically distributed (iid). However, in many cases involving the mining of text, voice and time series data, the data in step $t$ depends on the data from the previous steps. This brings us to sequential learning algorithms. Recurrent neural networks (RNN) are a class of machine learning algorithms used to treat data correlated in time as time series. However, RNN have a high computational cost, which makes them difficult to apply in the sequential learning scenario [15, 5].
[…]
Furthermore, these general predictive models do not provide us with a measure of confidence in the forecasts they make. In various applications, such as demand forecasting, it is important to establish a measure of confidence in the forecasts so that we can develop strategies for stock control. One strategy for developing a measure of confidence in forecasts is to use Bayesian techniques to quantify the uncertainty related to the model and estimates [13, 3]. Another alternative is to use “Conformal Prediction” developed by [16], which uses the assumption of independent observations to construct a set in the space of labels with a given probability [3]. Although the conformal prediction technique is distribution-free, it assumes that the data is independent.
The main objective of this work is to develop a class of algorithms that works incrementally that is suitable for large volumes of data updated online and has a time-dependent structure. In addition, we will propose a measure of confidence for the predictions made by the algorithm. To do this, we develop an extension of the class of SINDy models proposed by [2], incorporating random coefficients that are estimated and updated online. In this case, we leverage online convex optimization techniques to sequentially update the coefficients and to obtain probabilistic coverage guarantees for the prediction estimate.
This paper is organized as follows: in section 2, we introduce the proposed algorithm, randomized SINDy, and demonstrates its PAC learning properties through rigorous analysis. In section 3, we present experimental results, starting with simple simulations to verify the effectiveness of the proposed approach, followed by applications on two real-world datasets. Finally, section 4 concludes with a discussion of the results and future research directions.
2. Randomized SINDy
We consider the problem of sequential learning by the following online convex optimization setting. At the initial step, we observe $x_{1}\in\mathbb{R}^{k}$. Then the learner takes an initial action $A_{1}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{1}(x_{1},B_{1}):=\mathbb{E}_{b_{1}}\left(1\!\!1_{B_{1}}\Delta(x_{1})\beta_{1}\right),\quad B_{1}\in\beta(\mathbb{R}^{m}),$
where $\Delta(x)\in\mathbb{R}^{m\times p}$, for every $x\in\mathbb{R}^{k}$, is a feature library and $\beta_{1}$ has a $p$-dimensional multivariate distribution with density $g_{b_{1}}$ parametrized by $b_{1}\in\Theta\subset\mathbb{R}^{u}$. We assume that $\Theta$ is a bounded convex and compact set in the Euclidean space. We denote by $D$ an upper bound on the diameter of $\Theta$:
$\parallel b_{1}-b_{2}\parallel\leq D,\quad\forall~b_{1},b_{2}\in\Theta.$
As a consequence, the action $A_{1}$ can be identified by the parameter $b_{1}\in\Theta$. Next, we observe $y_{1}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{1},(x_{1},y_{1}))=\mathbb{E}_{b_{1}}\left(\ell(y_{1},\Delta(x_{1})\beta_{1})+\lambda\beta_{1}^{T}\beta_{1}\right)$
where $\lambda\geq 0$ is the complexity parameter and $\ell:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ is an integrable function. At the second step, we observe $x_{2}\in\mathbb{R}^{k}$. Then the learner takes an action $A_{2}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{2}(x_{2},B_{2}):=\mathbb{E}_{b_{2}}\left(1\!\!1_{B_{2}}\Delta(x_{2})\beta_{2}\right),\quad B_{2}\in\beta(\mathbb{R}^{m}),$
where $\beta_{2}$ has a $p$-dimensional multivariate distribution with density $g_{b_{2}}$ parametrized by $b_{2}\in\Theta$. As a consequence, the action $A_{2}$ can be identified by the parameter $b_{2}\in\Theta$. Next, we observe $y_{2}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{2},(x_{2},y_{2}))=\mathbb{E}_{b_{2}}\left(\ell(y_{2},\Delta(x_{2})\beta_{2})+\lambda\beta_{2}^{T}\beta_{2}\right).$
Following in this way, at the step $t$, we observe $x_{t}\in\mathbb{R}^{k}$. Then the learner takes an action $A_{t}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{t}(x_{t},B_{t}):=\mathbb{E}_{b_{t}}\left(1\!\!1_{B_{t}}\Delta(x_{t})\beta_{t}\right),\quad B_{t}\in\beta(\mathbb{R}^{m}),$
where $\beta_{t}$ has a $p$-dimensional multivariate distribution with density $g_{b_{t}}$ parametrized by $b_{t}\in\Theta$. As a consequence, the action $A_{t}$ can be identified by the parameter $b_{t}\in\Theta$. Next, we observe $y_{t}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{t},(x_{t},y_{t}))=\mathbb{E}_{b_{t}}\left(\ell(y_{t},\Delta(x_{t})\beta_{t})+\lambda\beta_{t}^{T}\beta_{t}\right).$
The goal of a sequential learning algorithm is to find a sequence $\{\hat{b}^{t}:t\geq 1\}$ that achieve the minimum regret along the whole learning process. The regret is defined as
$Regret=\sum_{t=1}^{n}L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-\min_{b\in\Theta}\sum_{t=1}^{n}L(b,(x_{t},y_{t})).$
In order to establish that the proximal gradient algorithm minimizes the regret, we suppose that the following hypothesis is valid.
Sequential convexity Hypothesis:
  • For every fixed $(x,y)\in\mathbb{R}^{k}\times\mathbb{R}^{m}$, the function $L(\cdot,(x,y)):\Theta\rightarrow\mathbb{R}$ is given by
    $L(b,(x,y))=\mathbb{E}_{b}\left(\ell(y,\Delta(x)\beta)+\lambda\beta^{T}\beta\right)$
    where $\beta$ has a $p$-dimensional multivariate distribution with density $g_{b}$ parametrized by $b\in\Theta$, is a convex and differentiable function.
  • $\parallel\nabla L(b,(x,y)\parallel\leq G$ with constant $G>0$.
As a consequence, we propose the following sequential proximal gradient algorithm
$\hat{b}^{t}({\bf o}_{t-1})=prox_{\eta_{t-1},\Psi}\left[\hat{b}^{t-1}({\bf o}_{t-2})-\eta_{t-1}\nabla L(\hat{b}^{t-1},(x_{t-1},y_{t-1}))\right]$
where $\hat{b}^{1}\in\Theta$, $t\geq 2$ and
$\operatorname*{prox}_{\eta,\Psi}(v)=\arg\min_{z}\left\{\Psi(z)+\frac{1}{2\eta}\parallel z-v\parallel^{2}\right\},$
with
$\Psi(b)=\begin{cases}0&;~~~b\in\Theta\\ \infty&;~~~b\not\in\Theta.\end{cases}$
Next, we apply usual arguments from the theory of convex learning to prove that the sequential proximal gradient minimizes the regret.
Theorem 2.1.
Under the sequential convexity hypothesis, the regret function satisfies
$\frac{1}{n}\left(\sum_{t=1}^{n}L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-\min_{b\in\Theta}\sum_{t=1}^{n}L(b,(x_{t},y_{t}))\right)\leq 3DG\frac{1}{\sqrt{n}},$
with $\eta_{t}=\frac{D}{G\sqrt{t}}$.
Proof.
We consider the following stochastic Ridge regression:
(2.1) $\hat{b}=\arg\min_{b\in\Theta}\mathbb{E}_{b}L(b,(x,y))=\arg\min_{b}\left\{L(b,(x,y))+\Psi(b)\right\}$
where $\lambda\geq 0$, $L(b,(x,y))=\mathbb{E}_{b}\ell(y,\Delta(x)\beta)+\lambda\mathbb{E}_{b}\beta^{T}\beta$ and
$\Psi(b)=\begin{cases}0&;~~~b\in\Theta\\ \infty&;~~~b\not\in\Theta.\end{cases}$
By convexity of $L$, we have that
(2.2) $L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))\leq L(b,(x_{t},y_{t}))+\langle\nabla L(\hat{b}^{t}({\bf o}_{t-1}),(x,y)),\hat{b}^{t}({\bf o}_{t-1})-b\rangle,\quad b\in\Theta.$
Applying the Pythagorean theorem, we have that
$\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}$ $=$ $\parallel prox_{\eta_{t},\Psi}\left[\hat{b}^{t}({\bf o}_{t-1})-\eta_{t}\nabla L(\hat{b}^{t},(x_{t},y_{t}))\right]-b\parallel^{2}$
$\leq$ $\parallel\hat{b}^{t}({\bf o}_{t-1})-\eta_{t}\nabla L(\hat{b}^{t},(x_{t},y_{t}))-b\parallel^{2}$
$=$ $\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}+\eta_{t}^{2}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}-2\eta_{t}\langle\nabla L(\hat{b}^{t},(x_{t},y_{t})),\hat{b}^{t}({\bf o}_{t-1})-b\rangle,$
for every $b\in\Theta$. Hence, we conclude that
(2.3) $2\langle\nabla L(\hat{b}^{t},(x_{t},y_{t})),\hat{b}^{t}({\bf o}_{t-1})-b\rangle\leq\frac{\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}-\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}}{\eta_{t}}+\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}.$
As a consequence of equations 2.2 and 2.3, we obtain that
$L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-L(b,(x_{t},y_{t}))$ $\leq$ $\frac{\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}-\parallel\hat{b}^{t+1}({\bf o}_{t})-b\parallel^{2}}{\eta_{t}}$
$+$ $\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}.$
Summing from $1$ to $n$ with $\eta_{t}=\frac{D}{G\sqrt{t}}$ and $\frac{1}{\eta_{0}}=0$, we obtain for every $b\in\Theta$ that
$\sum_{t=1}^{n}\left(L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-L(b,(x_{t},y_{t}))\right)$ $\leq$ $\sum_{t=1}^{n}\parallel\hat{b}^{t}({\bf o}_{t-1})-b\parallel^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)$
$+$ $\sum_{t=1}^{n}\eta_{t}\parallel\nabla L(\hat{b}^{t},(x_{t},y_{t}))\parallel^{2}$
$\leq$ $D^{2}\sum_{t=1}^{n}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{n}\eta_{t}$
$\leq$ $D^{2}\frac{1}{\eta_{n}}+DG\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\leq DG\sqrt{n}+DG2\sqrt{n}\leq 3DG\sqrt{n}.$
The last inequality follows since $\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\leq 2\sqrt{n}$.
If we consider the $p$-dimensional vector $\hat{\beta}_{t}$ with distribution $g_{\hat{b}_{t}}$, we can predict the output of the process as follows
$\hat{y}_{t}=\mathbb{E}_{\hat{b}_{t}}\left(\Delta(x_{t})\hat{\beta}_{t}\right),\quad t\geq 1.$
As we know the distribution of the quantity $\Delta(x_{t})\hat{\beta}_{t}$, the prediction probability interval is given by
$\mathbb{P}\left[LI_{\alpha/2}\leq\Delta(x_{t})\hat{\beta}_{t}\leq LS_{\alpha/2}\right]=1-\alpha,\quad 0<\alpha<1.$
2.1. Randomized SINDy for regression
Considering a sequential regression model $y=\Delta(x)\beta$ where $\beta=\left(\beta_{1},\cdots,\beta_{p}\right)$. If we assume $\beta$ follows a multivariate normal distribution $\mathcal{N}_{p}(\mu,\Sigma)$ with mean vector $\mu=(\mu_{1},\cdots,\mu_{p})$, covariance matrix $\Sigma=\operatorname{diag}(\sigma_{1},\cdots,\sigma_{p})$ and density function given by:
$g_{b}(\beta)=\frac{1}{(2\pi)^{p/2}|\Sigma|^{1/2}}\exp\left\{-\frac{1}{2}(\beta-\mu)^{\top}\Sigma^{-1}(\beta-\mu)\right\}.$
In this case, we have $b=\left(\mu,\Sigma\right)$, where $\mu\in\mathbb{R}^{p}$ and $\Sigma\in\mathbb{R}^{p\times p}$ (p.s.d.). By assuming this parametric form for $g$, the constraints for a probability density are inherently satisfied. At every time step, the density function $g_{b}(\beta)$ is completely determined by $b$. By using the quadratic loss, the optimization problem, according to Equation 2.1, is as follows:
(2.4) $\operatorname*{arg\,min}_{\mu,\Sigma}\left\{~\mathbb{E}~(y-\Delta(x)\beta)^{2}+\lambda\mathbb{E}~\beta^{\top}\beta+\Psi(\Sigma)~\right\}$
where $\lambda\geq 0$ and
$\Psi(\Sigma)=\begin{cases}0&;~\text{if }\Sigma\text{ is positive semi-definite (p.s.d.),}\\ \infty&;~\text{otherwise}.\end{cases}$
We now compute each term in the objective function 2.4. Let $z:=\Delta(x)$ and since $\beta_{p}\sim\mathcal{N}(\mu,\Sigma)$,
$(y-z^{\top}\beta)^{2}=y^{2}-2yz^{\top}\beta+(z^{\top}\beta)^{2},$
the expected quadratic loss becomes
$\mathbb{E}(y-z^{\top}\beta)^{2}$ $=$ $\mathbb{E}y^{2}-2yz^{\top}\mathbb{E}\beta+\mathbb{E}(z^{\top}\beta)^{2}.$
We know that
$\mathbb{E}[\beta]=\mu,$
$\text{Var}(\beta)=\Sigma=\mathbb{E}[\beta\beta^{\top}]-\mathbb{E}[\beta]\mathbb{E}[\beta]^{\top}\implies\mathbb{E}[\beta\beta^{\top}]=\Sigma+\mu\mu^{\top}.$
Thus,
$\mathbb{E}~(z^{\top}\beta)^{2}$ $=$ $\mathbb{E}~z^{\top}\beta\beta^{\top}z$
$=$ $z^{\top}\mathbb{E}[\beta\beta^{\top}]z$
$=$ $z^{\top}(\Sigma+\mu\mu^{\top})z$
$=$ $z^{\top}\Sigma z+z^{\top}\mu\mu^{\top}z$
$=$ $z^{\top}\Sigma z+(z^{\top}\mu)^{2}.$
Substituting these into the expected quadratic loss, we obtain
$\mathbb{E}\left[(y-z^{\top}\beta)^{2}\right]=y^{2}-2yz^{\top}\mu+z^{\top}\Sigma z+(z^{\top}\mu)^{2}.$
For $\lambda>0$, the expectation of $L_{2}$ regularization term is simply
$\lambda\mathbb{E}\beta^{\top}\beta=\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right).$
And the constraint term enforces that the estimated covariance matrix of $\beta$ is p.s.d. which can be represented using an indicator function:
$1\!\!1_{p.s.d}:\Sigma\rightarrow\{0,1\}.$
We can apply proximal gradient descent to minimize the objective function which can be treated as two parts, $f_{1}$ and $f_{2}$:
$\underbrace{y^{2}-2yz^{\top}\mu+z^{\top}\Sigma\mathbf{x}+(z^{\top}\mu)^{2}+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
The partial derivative with respect to $\mu$ is:
$\nabla_{\mu}f_{1}(\mu,\Sigma)=2z(z^{\top}\mu-y)+2\lambda\mu,$
and the partial derivative with respect to the $\Sigma$ is:
$\nabla_{\Sigma}f_{1}(\mu,\Sigma)=zz^{\top}+\lambda I.$
The indicator function in $f_{2}$ is non-smooth and the proximal operator is reduced to a projection onto a p.s.d. cone, following [1],
$\text{proj}_{K}(\Sigma)=\sum_{i=1}^{p}\max\{0,\zeta_{i}\}v_{i}v_{i}^{\top},$
where $K$ is the p.s.d. cone and $\Sigma=\sum_{i=1}^{p}\zeta_{i}v_{i}v_{i}^{\top}$ is the eigenvalue decomposition of $\Sigma$.
Therefore, at the end of every time step, we update the parameters of $g_{t+1}$ based on the observed data $(z_{t},y_{t})$:
(2.5) $\mu^{t+1}=\mu^{t}-\eta_{t}\nabla_{\mu^{t}}f_{1}(\mu^{t},\Sigma^{t})=\mu^{t}-\eta_{t}\left(2z_{t}\left(z_{t}^{\top}\mu^{t}-y_{t}\right)+2\lambda\mu^{t}\right),$
$\widetilde{\Sigma}^{t+1}=\Sigma^{t}-\eta_{t}\nabla_{\Sigma^{t}}f_{1}(\mu^{t},\Sigma^{t})=\Sigma^{t}-\eta_{t}\left(z_{t}z^{\top}_{t}+\lambda I\right),$
(2.6) $\Sigma^{t+1}=\text{proj}_{K}\left(\widetilde{\Sigma}^{t+1}\right)$
where $\eta_{t}>0$, $\lambda\geq 0$.
The prediction for the next output at step $t+1$, given the observed feature vector $z_{t+1}$, is computed as
$\hat{y}_{t+1}=z_{t+1}^{\top}\hat{\beta}_{t},\text{ where }\hat{\beta}_{t}=\mu^{t}$
and the corresponding $100(1-\alpha)\%$ prediction probability interval is given by
$\hat{y}_{t+1}\pm z_{\alpha/2}\sqrt{\widehat{Var}(\beta)_{t}},\text{ where }\widehat{Var}(\beta)_{t}=z_{t+1}^{\top}\Sigma^{t}z_{t+1}\text{ and }0<\alpha<1.$
2.2. Randomized SINDy for binary outcome
In a binary classification setting where $Y\in\{0,1\}$ and $z:=\Delta(x)\in\mathbb{R}^{p}$, we can adapt the previous example by utilizing the logistic function such that $h(z):=\sigma(z^{\top}\beta)$ represents the probability of predicting the output label $Y=1$:
$P(Y=1\mid z,\beta)=\sigma(z^{\top}\beta)=\dfrac{1}{1+\exp\{-z^{\top}\beta\}}.$
Using the logistic loss function,
$\ell(y,z^{\top}\beta)=-\bigr[y\ln\left(h\left(z\right)\right)+\left(1-y\right)\ln\left(1-h\left(z\right)\right)\bigr],$
we note that there is no closed form for $\mathbb{E}\ell(y,z^{\top}\beta)$. Rather, we consider its approximation $\ell(y,z^{\top}\mu)$. Consequently, the objective function is as follows:
$\underbrace{\ell(y,z^{\top}\mu)+\lambda\left(\text{tr}\Sigma+\mu^{\top}\mu\right)}_{f_{1}}+\underbrace{1\!\!1_{p.s.d}(\Sigma)}_{f_{2}}$
The partial derivative with respect to $\mu$ is:
$\nabla_{\mu}f_{1}(\mu,\Sigma)=z\left(\sigma\left(z^{\top}\mu\right)-y\right)+2\lambda\mu,$
and the partial derivative with respect to the $\Sigma$ is:
$\nabla_{\Sigma}f_{1}(\mu,\Sigma)=\lambda I$
Therefore, at the end of every time step, we update the parameters of $g_{t}$ based on the observed data $(\mathbf{x}_{t},y_{t})$:
$\mu_{t+1}=\mu_{t}-\eta_{t}\nabla_{\mu_{t}}f_{1}(\mu_{t},\Sigma_{t})=\mu_{t}-\eta_{t}\left(\mathbf{x}_{t}\left(\frac{1}{1+e^{-z^{\top}_{t}\mu_{t}}}-y_{t}\right)+2\lambda\mu_{t}\right),$
$\widetilde{\Sigma}_{t+1}=\Sigma_{t}-\eta_{t}\nabla_{\Sigma_{t}}f_{1}(\mu_{t},\Sigma_{t})=\Sigma_{t}-\eta_{t}\left(\lambda I\right),$
$\Sigma_{t+1}=\text{proj}_{K}\left(\widetilde{\Sigma}_{t+1}\right)$
where $\eta_{t}>0$, $\lambda\geq 0$.
3. Experiments
In this section, we present some simulation results of models discussed in section 2. All experiments were executed using the R language.
Experiment 1
Considering an outcome sequence $Y=2X+\varepsilon$ where $\varepsilon\sim N(0,\epsilon)$, the data is generated by selecting $X\sim U(-2,2)$. To analyze the impact of noise and the choice of learning rate $\eta$, we vary the noise level between 0.1 and 1. For the low-noise case ($\epsilon=0.1$), we consider learning rates of 0.01 and 0.1. In the high-noise case ($\epsilon=1$), we vary $\eta$ between 0.03 and 0.3. In both cases, the initial value for parameter is set to $\mu_{0}=0$, also, no regularized term is considered ($\lambda=0$). We simulated a sequence of size 120 and the model quality results are summarized in Table 1. Additionally, Figure 1 presents a comparison between the true values of $y$ and the predicted values across different scenarios.
Table 1. Simulation results
$\epsilon$ $\eta$ $R^{2}$ $\hat{\sigma}$ RSME
0.1 0.01 0.8490 0,9305 0,9267
0.1 0.9805 0.3341 0.3327
1 0.03 0.7973 1.1933 1.188
0.3 0.4156 2.0260 2.0175
Figure 1. Study of the impact of noise on the selection of learning rate
For low-noise data, a higher learning rate achieves a better R-squared value and an estimated $\sigma$ closer to the true noise level. However, in high-noise scenarios, a lower learning rate is preferable, as it helps stabilize learning and prevent erratic fluctuations in parameter updates. As observed in the graph, higher learning rate results in larger prediction errors. This suggests that a higher learning rate prioritizes adjustments based on recent data, while a lower learning rate means to retain information given by older data entries. We conducted a Monte Carlo simulation with 1,000 replicates, each consisting of a sequence size of 10,000 observations. The simulation results are shown in Table 2.
Table 2. Simulation results
$\epsilon$ $\eta$ $R^{2}$ $\hat{\sigma}$ RSME $\hat{\beta}$
0.1 0.1 0.9976 0.1144 0.1144 2.000 (1.9979, 2.002)
1 0.03 0.8346 1.0232 1.0232 2.004 (1.9893, 2.0107)
To evaluate the impact of the initial value, we vary $\mu_{0}$ between 0 and 1.8. As shown in Figure 2, the initial values influence the convergence rate. For a sequence observed at the 50th time step, the initial value of $\mu_{0}=0$ requires more iterations to achieve convergence.
Figure 2. Study of the impact of initial values on convergence
Experiment 2
We created an ill-conditioned model matrix for simulation purpose and the predictors were generated as follows: $X_{1}\sim\mathcal{N}(0,1)$, $X_{2}=X_{1}+\delta_{2}$, $X_{3}=2X_{1}+3X_{2}+\delta_{3}$, where $\delta_{2}\sim N(0,0.01^{2})$, $\delta_{3}\sim N(0,0.01^{2})$. The outcome sequence was defined as $Y=-1+2X_{1}-2X_{2}+1.5X_{3}+\varepsilon$, with $\varepsilon\sim N(0,1)$. A sequence of 500 time steps was generated.
The first 250 time steps were designated as a warm-up period for model estimation. Within this warm-up period, we performed a rolling window cross-validation with 40 train-validation splits. The first training set consisted of 50 observations. For each subsequent split, the training set size was incremented by five observations (e.g.. 50, 55, …, up to 245 observations). Each corresponding validation set contained the five subsequent observations immediately following its training set. For each of these 40 train-validation splits, we fitted a regularized regression model. The regularization parameter ($\lambda$) was varied across 30 different values, ranging from 0 to 1. To determine the optimal $\lambda$, the minimum Root Mean Squared Error (RMSE) was calculated for each $\lambda$ value on each of the 40 validation sets. Subsequently, $\lambda=0.3448$ was selected as the optimal $\lambda$ as it yielded the lowest overall RMSE. Still within this warm-up period, we proceeded to standardize the model matrix, which has dimensions of $250\times 4$. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0})$. Subsequently, the algorithm was applied with the parameter update steps defined by Equations 2.5 and 2.6. The model parameters estimated at $t=250$ were used as the initial state for the subsequent time step. From $t=251$ onwards, each new incoming data dynamically updated both the standardization of the model matrix and the estimated parameters. With a learning rate of $\eta=1\times 10^{-3}$, Table 3 presents the estimated parameters at the initial and the final time step.
Table 3. Estimated parameters at initial and final iterative step of the warm-up estimation period
Parameter $t$ intercept $X_{1}$ $X_{2}$ $X_{3}$
$\hat{\mu}_{t}$ 1 -1.5652 2.5442 2.2182 2.2142
500 -1.3349 2.2946 2.0639 2.0610
$\widehat{\Sigma}_{t}$ 1 0.0075 0.1271 0.0952 0.0112
500 0.0000 0.0267 0.0195 0.0006
Model quality metrics at the final time step are presented in Table 4:
Table 4. Model quality measure at the final time step
$t$ SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
500 25216.07 738.6398 500 4 0.97071 1.22032 1.21543
Experiment 3
We simulated a binary sequence $Y\sim Ber(p)$, where $p$ was defined by the logistic model $P(Y=1\mid X)=\dfrac{1}{1+\exp\{-(1+2X_{1}+3X_{2})\}}.$ As in the previous experiment, a sequence of 500 time steps was generated. Initial logistic model parameters were obtained by fitting a ridge logistic regression with $\lambda=0$. The algorithm from Example 2 of section 2 was then applied with a fixed learning rate of $\eta=0.1$ . Table 5 shows the estimated coefficients at both the initial and final time steps.
Table 5. Estimated parameters at initial and final iterative step of the warm-up estimation period
$t$ intercept $X_{1}$ $X_{2}$
1 0.7897 1.9020 2.8405
500 0.8002 1.9081 2.8414
Figure 3. ROC curve
The Receiver Operating Characteristic (ROC) curve, Figure 3, indicated an optimal threshold of 0.5745, resulting in an Area Under the Curve (AUC) of 0.925. Additional model quality metrics, such as an accuracy greater than 80% (Table 6), confirm the model’s strong performance.
Table 6. Model quality measure
Accuracy TPR TNR Precision F1 Entropy
0.842 0.8522 0.8278 0.8732 0.8626 2.5467
Experiment 4
We applied the models discussed in section 2 to forecast the monthly U.S. unemployment rate. As described in the work of [8], the independent variables included are Initial Claims (IC) for unemployment insurance, the inflation rate (measured as the seasonal growth rate of the Consumer Price Index), and industrial growth (the seasonal growth rate of the Industrial Production Index). All data were sourced from the FRED database (https://fred.stlouisfed.org/). The analysis period spans from January 1967 to May 2025. Table 7 gives a brief description of the variables used for forecasting. While most variables are originally monthly, the weekly Initial Claims data was averaged to a monthly frequency.
Table 7. Variable descriptions for U.S. unemployment rate forecasting, data covering January 1967 to May 2025
Variable Description
UNEMP The monthly unemployment rate (as a percentage)
UNEMP_LAG1 The monthly unemployment rate (as a percentage) from previous month
IC The number of Initial Claims (IC) of unemployment insurance
CPI The Consumer Price Index (CPI)
IPI The Industrial Production Index (IPI)
Since the unemployment rate is non-negative, we applied a transformation to the response variable $Y=\log\left(\frac{UNEMP}{100-UNEMP}\right)$. Based on the Akaike Information Criterion (AIC), the inclusion of the variables in Table 7 is validated using the auto.arima function from the forecast package. The model that included the lagged unemployment rate and three other explanatory variables obtained the smallest AIC as shown in Table 8.
Table 8. AIC comparison for forecasting models
Model AIC
UNEMP_ LAG1 -1661.203
UNEMP_ LAG1 & IC -1700.652
UNEMP_ LAG1, IC & CPI -1761.693
UNEMP_ LAG1, IC, CPI & IPI -1772.673
Figure 4. Train-validation set splits for hyperparameter selection
We used the initial 30 years of data, spanning from January 1967 to December 1999, as a warm-up period for our online learning model. Based on a rolling expanding window approach, we began with 20 years of data for training to make a one-year-ahead prediction. This process generated 12 distinct train-validation set splits for model evaluation. For each of these 12 training sets, we performed regression with Tikhonov regularization, exploring 30 different regularization parameter ($\lambda$) values ranging from 0 to 1. The optimal regularization term was identified as being the value of 0, as this produced the minimum overall RMSE (see Figure 4). We then proceeded to standardize the model matrix, which had dimensions of 395 × 5. A preliminary Tikhonov regularization was then performed to obtain an initial guess for the mean coefficient vector ($\mu_{0}$) and its covariance matrix ($\Sigma_{0}$).The algorithm was then applied with the parameter update steps defined by Equations 2.5 and 2.6. Table 9 presents the estimated parameters from both the initial guess and the final iterative step of the warm-up estimation period.
Table 9. Estimated parameters at initial and final iterative step of the warm-up estimation period
Parameter $t$ intercept UNEMP_LAG1 IC CPI IPI
$\hat{\mu}_{t}$ 1 -2.7536 0.2461 0.0272 0.0054 -0.0132
395 -2.7533 0.2454 0.0267 0.0051 -0.0137
$\widehat{\Sigma}_{t}$ 1 0.0064 0.0227 0.0204 0.0986 0.0968
395 0.0051 0.0079 0.0086 0.0971 0.0835
Figure 5 illustrates the predicted employment rates along with the true employment rates on the original scale. The close alignment between the predicted and actual values is evident, with an $R^{2}$ of 0.99 indicating a very strong fit (Table 10). Using the estimated parameters from the final warm-up period time step ($t=395$), we made a forecast for the following month’s unemployment rate. Table 11 presents this forecasted value with its 95% uncertainty level, which can be compared to the true observed value in both transformed and original scale (in brackets). The results of the residual analysis are presented in Figure 6.
Figure 5. U.S. unemployment rate spanning from January 1967 to January 2000. Predicted values (blue), actual values (gray) and January 2000 forecast (red dot)
Table 10. Model quality measure at the final time step in warm-up estimation period
Date SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
1999-12 29.1597 0.2716 395 5 0.9907 0.0264 0.0262
Table 11. Forecasted value and prediction interval
Date Forecasted Lower limit Upper limit True value
2000-01 -3.1836 -3.3108 -3.0563 -3.1781
(3.9789) (3.5201) (4.4947) (4)
Values
Upper Limit 0.0821
Central Line 0.0004
Lower Limit -0.0813
(a) Individual residual values chart
(b) Observed v.s. fitted values
(c) Residuals plot
Values
Min. -0.0851
1st Qu. -0.0165
Median -0.0006
Mean 0.0004
3rd Qu. 0.0159
Max. 0.0912
(d) Exploratory of residuals
Figure 6. Residuals analysis for the warm-up estimation period
The Individual values control chart in Figure 6 confirms that the residuals are randomly distributed with a zero mean. This interpretation is consistent with the residuals plot in 6(c). Additionally, the histogram in Figure 6 displays a symmetric distribution of residuals centered around zero.
Utilizing the final estimated model from the warm-up period, we then proceeded to enter a real-time estimation scenario. From January 2000 to May 2025, data were processed incrementally. At every time step, each new incoming data dynamically updated both the standardization of the model matrix and the estimated parameters. Figure 7 illustrates that, in general, the forecasts are close to the true unemployment rates throughout the entire period. Two specific intervals were noted as exceptions. The first deviation began in 2009, coinciding with the Great Recession. The second deviation began in the first quarter of 2020, precisely corresponding with the onset of the COVID-19 pandemic. The results of the residual analysis are presented in Figure 8.
Figure 7. U.S. unemployment rate spanning from January 2000 to May 2025. Predicted values (blue) and actual values (gray)
Analysis of the individual values control chart for residuals (Figure 8) confirms that the observed prediction deviations during the COVID-19 pandemic are indeed outliers. A complete list of all the outliers identified (points that failed the first Nelson test) can be found in Table 12. We can see from the same figure that, starting from somewhere in 2008, the individual residual values present an increase trend, which indicates that the algorithm underestimated the unemployment rate. After a certain amount of time, the algorithm adapted to the underlying data drift, as evidenced by the gradual convergence of residuals to the central zero line. However, these slight deviations were not identified as outliers. The second deviation period, coinciding with the COVID-19 pandemic, caused a massive shock, with the unemployment rate spiking by over 10% , from 4.4% in March 2020 to 14.8% in May 2020. Due to this high shock, the algorithm produced predictions that were slightly higher than usual. These two observations clearly demonstrate the adaptive capabilities of our algorithm to sudden or abrupt changes in the data. Table 13 shows the changes in coefficient values for these outlier data points.
Values
Upper Limit 0.0954
Central Line -0.0130
Lower Limit -0.1213
(a) Individual residual values chart
(b) Observed v.s. fitted values
(c) Residuals plot
Values
Min. -0.5138
1st Qu. -0.0319
Median -0.0085
Mean -0.013
3rd Qu. 0.0159
Max. 0.6357
(d) Exploratory of residuals
Figure 8. Residuals analysis for the real-time learning scenario
Table 12. Outliers detected by Nelson test 1 in individual residual values chart
Date Forecast True value Lower limit Upper limit
2020-03 4.4 6.9266 0.9807 35.8646
2020-04 14.8 8.4239 5.4525 12.7955
2020-05 13.2 20.2683 14.4422 27.6841
2020-06 11.0 14.7956 12.3572 17.6185
2020-07 10.2 12.0806 10.4351 13.9451
2020-08 8.4 10.3500 8.8846 12.0251
2020-10 6.9 8.0244 7.3180 8.7925
Table 13. Changes of coefficients of outlier data point
Date intercept UNEMP_LAG1 IC CPI IPI
2020-02 -2.7531 0.2524 0.0281 0.0056 -0.0140
2020-03 -2.7537 0.2536 0.0173 0.0045 -0.0147
2020-04 -2.7519 0.2517 0.0550 0.0075 -0.0136
2020-05 -2.7528 0.2483 0.0471 0.0059 -0.0143
2020-06 -2.7535 0.2463 0.0438 0.0049 -0.0148
2020-07 -2.7539 0.2454 0.0421 0.0042 -0.0152
2020-09 -2.7545 0.2442 0.0406 0.0031 -0.0160
The application of the first Nelson test for outlier detection suggests its efficacy in identifying early concept drift in online learning algorithms. To visualize this, Figure 9 presents the evolution of $\hat{\sigma}$ and RMSE over 305 time steps. Visually, the onset of concept drift consistently aligns with the position of the first detected outlier. Detailed metrics for outlier data points are presented in Table 14.
Figure 9. Metric monitoring
Table 14. Model quality measure at selected time steps in real-time learning
Date SSTSST SSESSE $n$ $p$ $R^{2}$ $\hat{\sigma}$ RMSERMSE
2020-03 53.0959 0.7343 638 5 0.9862 0.0341 0.0339
2020-04 54.1529 1.1384 639 5 0.9790 0.0424 0.0422
2020-05 54.9513 1.4024 640 5 0.9745 0.0470 0.0468
2020-06 55.4205 1.5180 641 5 0.9726 0.0488 0.0487
2020-07 55.7799 1.5542 642 5 0.9721 0.0494 0.0492
2020-08 55.9279 1.6073 643 5 0.9713 0.0502 0.0500
2020-10 56.0493 1.6463 645 5 0.9706 0.0507 0.0505
2025-05 64.7679 1.8334 700 5 0.9717 0.0514 0.0512
Table 15 shows the estimated parameter values from both the initial and final iterative steps. The learning rate used was set to $\eta=1\times 10^{-3}$ across all iterations. For each of the 305 time steps in the real-time forecast (from January 2000 to May 2025), we calculated a 95% prediction level using the estimated covariance matrices of the coefficients. The empirical coverage of this prediction interval was found to be 0,9639. Out of these 305 prediction intervals, 11 failed to contain the true unemployment rates.
Table 15. Comparison of estimated parameters at the initial and final real-time learning iterative step (Feb 1967 and May 2025)
Parameter $t$ intercept UNEMP_LAG1 IC CPI IPI
$\hat{\mu}_{t}$ 1 -2.7536 0.2461 0.0272 0.0054 -0.0132
700 -2.7603 0.2495 0.0411 -0.0081 -0.0232
$\widehat{\Sigma}_{t}$ 1 0.0064 0.0227 0.0204 0.0986 0.0968
700 0.0051 0.0008 0.0018 0.0535 0.1299
Experiment 5
In this experiment, we applied the randomized SINDy algorithm to the Elec2 dataset, a real-world binary classification problem. The objective of the dataset focuses on predicting the direction of electricity price changes (up or down) in New South Wales (NSW), Australia. Comprising 45,312 instances, the dataset includes five features and spans the period from May 7, 1996 to December 5, 1998. Table 16 provides a brief description of the variables and further details can be found in [7]. The dataset was sourced from Kaggle (https://www.kaggle.com/datasets/yashsharan/the-elec2-dataset). All features, with the exception of “day”, were pre-normalized to the range of 0 and 1 upon download. To maintain a uniform data scale, the “day” variable was subsequently normalized using min-max scaling. The class label distribution is shown in Table 17.
Table 16. Variable descriptions for Elec2, data covering
Variable Description
class NSW price higher (UP-1) or lower (Down-0) than 24 hour average
day Day of week
period Time (based on 30-minute periods)
nswdemand NSW electricity demand
vicdemand Victoria electricity demand
transfer Scheduled electricity transfer between states
Table 17. Class label distribution
Up (1) Down (0)
Count 19237 26075
Proportion 0.5755% 0.4245%
We utilized the initial ten thousand instances for warm-up estimation in order to determine the hyperparameter and the initial guess values for the weights in the logistic function. This procedure involved 25 train-validation set splits. The first split began with 1344 instances in the training set used to do a one-week-ahead prediction corresponding to $7\times 48=336$ instances. The optimal regularization term, $\lambda=0.8276$, was chosen based on the highest accuracy achieved across all 25 validation sets. Subsequently, a regularized logistic regression was pre-run with the chosen lambda to obtain the initial weight values for the logistic function. With these initial weight values, we then applied the proposed algorithm, updating the weights at each time step. The estimated parameters of both the initial and final iterative steps are presented in Table 18.
Table 18. Estimated parameters at initial and final iterative step of the warm-up estimation period
Parameter $t$ intercept day period nswdemand vicdemand transfer
$\hat{\mu}_{t}$ 1 -0.0493 0.0912 0.0426 0.7496 0.1292 0.1267
45312 -0.3103 -0.3103 -0.3113 -0.0876 -0.1101 -0.074
Finally, the classification threshold was set to 0.5654, which was derived from the ROC curve (AUC 0.8313), in 10(a). The accuracy and other performance metrics are presented in 10(b). The ROC curve shows that our model has strong predictive power with high accuracy (above 80%).
(a) ROC curve
Metric Values
Accuracy 0.8533
TPR 0.8272
TNR 0.8726
Precision 0.8273
F1 0.8272
Entropy 1.5377
(b) Metrics
Figure 10. Model performance
The Elec2 dataset is widely recognized as a benchmark for concept drift detection in the online learning literature. Following the Drift Detection Method (DDM) proposed by [6], the trace of model’s error rate, $p_{i}$, the probability of a false prediction up to step $i$), is presented in Figure 11. The left panel of Figure 11 shows no apparent major changes in the overall error rate. The right panel, a zoomed-in view focusing on the range of 0.12 to 0.16, reveals a slight increase in the error rate within an early window. However, after this period, the error rate consistently decreases and then stabilize, fluctuating around a level of 0.146 indicated by the red dashed line. This observed error trace differs from those presented in the study by [6] for the same dataset, which shows various moments of abrupt drifts. This discrepancy may be attributed to the dynamic and adaptive learning capabilities of our proposed model, highlighting its ability to adjust to changes in the data distribution even when resources for explicit major drift alerts are unavailable.
Figure 11. Error rate tracing: y-axis spanning the full range from 0 and 1 (left); y-axis zoomed into the range from 0.12 to 0.16(right)
4. Discussion and Conclusion
In this paper, we propose a randomized sequential machine learning algorithm capable of handling dynamic data structures that does not rely on the classical iid assumption. Section 2 provides an introduction to the dynamic data arrival structure from which the algorithm was developed from a probabilistic perspective. We also provide in this section an extensive proof that mathematically demonstrates the proposed algorithm satisfies the PAC learning property. Additionally, a proof-of-concept was presented in the form of a theorem and a related proof. The proposed randomized algorithm considers a class of hypotheses or predictors that fulfill a probability density function. The learner takes action dynamically at every time step according to this density. More precisely, the learner uses the probability distribution of the predictors learned from the previous time step to make a prediction for the current time step and updates the density after the incoming data for the current time step is revealed. At each time step, the algorithm updates the weights assigned to the predictors using a gradient descent and a proximal algorithm, the latter being which is equivalent to a projection operator and ensures that the distributed weights form a valid probability density.
We were inspired by the SINDy (Sparse Identification of Nonlinear Dynamics) algorithm, a widely used data-driven modeling method for dynamics systems in recent years. We named our proposed algorithm “randomized SINDy” which considers the possibility of augmenting independent variables. This augmentation is sometimes referred to as dictionary or feature library, resulting in a high-dimensional model matrix that admits a sparse representation. We introduced Tikhonov regularization to the objective function to overcome the problem of ill-conditioned data. Assuming the weights follow a multivariate normal distribution, the proximal step is unnecessary. Thus, the learner’s task is to estimate the parameters of the multivariate normal distribution. We explored two common applications, regression problem and binary classification with a logistic model, both of which include a regularized term in the objective function.
Section 3 presents five experiments using simulated data and application on real-world data. The results of these experiments indicate the usefulness of our proposed algorithm.
In conclusion, we present some observations and ideas for future studies, which are listed below.
4.1. A smarter way to estimate an optimal learning rate and strategies for initial guess
The learning rate was determined through empirical testing during our experimentation. It is essential to seek a more intelligent alternative to identify an optimal learning rate and hyperparameter. Additionally, the convergence rate of the algorithm depends on the initial value. How should a strategy be defined to obtain these initial guesses? How can a reasonable value be determined for the regularized term? One approach involves using a rolling or expanding window to obtain the train-validation set splits. For each set, the regression or classifier is applied to the training set and predictions are made on the corresponding validation set. Also, we vary the range of values for $\lambda$ in each model fitting. The optimal $\lambda$ is identified as the one that produces the most optimal overall performance metrics across the validation sets. However, there are challenging issues using this approach, such as the determination of the window size and the quantity of available data for warm-up estimation.
4.2. Normalization procedures
We applied standardization to model matrix in our experiments. A disadvantage of this normalization process is its data dependency. In real-time settings, data is dynamically fed, necessitating continual updates to the mean and standard deviation statistics. Alternative normalization procedures, such as Min-Max scaling or data-independent functions like tan and arc-tan, could be explored.
4.3. Seeking for a better approach to derive the update step for covariance matrix in classification
The logistic model parameter update steps in Example 2 of Section 4 simplify the estimation procedure, as there is no exact derivative form. These simplifications for computational tractability may underestimate variation of the coefficients.
4.4. Computational cost
The proposed algorithm aims to provide a scalable structure that can efficiently handle large volumes of data. Therefore, it is important to conduct a simulation study and provide an estimation of the computation cost in terms of time consumption and memory usage.
4.5. Performance comparisons
A further study can be conducted by comparing the performance of the proposed algorithm on real-world data with that of some commonly applied classical methods.
4.6. Strategies on concept drift detection
As we saw in experiment 5, the individual residual values chart served as a useful tool for detecting potential data drift. Other Nelson tests could also be employed to identify anomalous trends in residuals, which may signal either data or concept drift. Further exploration into the capabilities of the proposed algorithm for detecting concept drift and formulating effective strategies is necessary.
4.7. Study expansion
To expand the study and include the case of infinite hypotheses and use related techniques to abstract Wiener space to build probability in the predictor space.